use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
z^2 + z + 1 = 0

Answer by venugopalramana(3286)

use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
z^2 + z + 1 = 0
D=B^2-4AC=1-4=-3
NO REAL ROOTS....2 COMPLEX CONJUGATE ROOTS